Block 1, of mass m1 = 0.700 kg , is connected over an ideal (massless and frictionless) pulley to block 2, of mass m2, as shown. For an angle of θ = 30.0 ∘ and a coefficient of kinetic friction between block 2 and the plane of μ = 0.300, an acceleration of magnitude a = 0.300 m/s2 is observed for block 2. Find mass of block 2
Answer #1

Tension, T = M2a + M2g sin(theta) + uM2g cos(theta)

T = M2(a + g sin(theta) + ug cos(theta))

T = m1 (g - a)

m1 (g - a) = M2(a + g sin(theta) + ug cos(theta))

M2 = m1 (g - a)/[a + g sin(theta) + ug cos(theta)]

M2 = [0.7 (9.8 - 0.3)]/[0.3 + (9.8 x sin(30) + (0.3 x 9.8 cos(30))]

M2 = 0.86 kg

Comment in case any doubt please rate my answer.....